OpenStudy (anonymous):

find the f-1(x) of f(x) =2x+1 over x-3

OpenStudy (anonymous):

\[y= \frac{ 2x+1}{x-3}\]

OpenStudy (anonymous):

swap x and y

OpenStudy (anonymous):

\[x = \frac{2y +1}{y-3}\]

OpenStudy (anonymous):

now x(y-3) = 2y+1 xy -3x = 2y +1 Now get the y's together and factor xy-2y = 1+3x y(x-2) = 1+3x y= (1+3x) / (x-2)

OpenStudy (anonymous):

f ^(-1) (x) = (1+3x) / (x-2)